OpenStudy (anonymous):

solve for x: 3|x-3|+2=14

mathslover (mathslover):

\[3|x-3|+2=14\] \[3(x-3)+2=14\] \[3x-9=14\] \[3x=23\] \[x=23/3\]

mathslover (mathslover):

we have taken |x-3|as positive

OpenStudy (lgbasallote):

@mathslover you seem to have omitted 2 ;) first subtract 2 from both sides 3|x+3| = 12 now divide both sides by 3 |x+3| = 4 now solve for x when: x + 3 = 4 AND x + 3 = -4 do you get it @catstar98 ?

OpenStudy (lgbasallote):

or do you need further instructions?

mathslover (mathslover):

sorry for that mistake